OpenStudy (anonymous):

Find least positive residue of 2^32 modulo 47

OpenStudy (anonymous):

Please show steps

OpenStudy (kc_kennylau):

2^1 % 47 = 2 2^2 % 47 = 4 2^4 % 47 = 16 2^8 % 47 = 16*16 % 47 = 21 2^16 % 47 = 21*21 % 47 = 18 2^32 % 47 = 18*18 % 47 = 42

OpenStudy (anonymous):

What are the steps you are taking. I see them but I am not exactly sure why you are taking them steps. For instance, why are are going to 2^8? Why are you choosing the exponents that you are choosing? Why 16*16/ or 21*21?

OpenStudy (kc_kennylau):

2^1 % 47 = 2 2^2 % 47 = 4 2^4 % 47 = 16 2^8 % 47 = (2^4 % 47)^2 % 47 = 16*16 % 47 = 21 2^16 % 47 = (2^8 % 47)^2 % 47 = 21*21 % 47 = 18 2^32 % 47 = (2^16 % 47)^2 % 47 = 18*18 % 47 = 42

OpenStudy (anonymous):

is it safe to say that you squared each time. (2^1)^2=(2^2), (2^2)^2= 2^4 and so on

OpenStudy (kc_kennylau):

Yes it's safe

OpenStudy (anonymous):

Great thank you